OpenStudy (anonymous):

What number must be added to the expression below to complete the square? x^2 - x A. 1/2 B. -1/4 C. -1/2 D. 1/4

OpenStudy (eric_d):

(a+b) ^ 2 = a^2 + b^2 + 2*a*b

OpenStudy (anonymous):

well,yes in this way u have 2ab but in that equation it's -x so u will discover that b=-1,2 ! so \(x^2 - x - \color{red}{\frac{1}{4}} \) make it complete square.

OpenStudy (eric_d):

@LewisSparrow96